An organic compound weighing 6g on heating with NaOH gave NH3 which is neutralised by 200mL of 1N of HCl. Percentage of nitrogen in it is


1) 12%


2)60%


3)46.67%


4)26.67%


I want the answer with detailed calculation.

Grade:12

1 Answers

Sunil Kumar FP
9 years ago
let the organic compund be represented by X
x +NaOH ------NH3 +HCL----NH4Cl
mole of NH3 formed is 200*1/1000=.2mole
weight of nitrogen in NH3=.2*14=2.8gm
therefore weight in NH3 and weight of nitrogen in organic compound will be same
thus the percentage of nitrogen in organic compound =(2.8/6)*100=46.67%